To graduate, Qyrisha needs at least a B in Statistics. She did not do very well on her first three tests. However, she did well in the last four. Below are her scores:

57 60 84 90 92 95

Compute the mean
Determine whether Qyrisha’s grade will be a B (70 to 79 average) or a C (60 to 69 average)
1
Expert's answer
2019-08-20T09:35:05-0400

Compute mean: (a1+a2+...+an)/n=(57+60+84+90+92+95)/6=79.67. Therefore Qyrisha’s grade will be a B.